# Antithetic Noise in Diffusion Models

### v1

@@ Abstract
We initiate a systematic study of antithetic initial noise in diffusion models. Across unconditional models trained on diverse datasets, text-conditioned latent-diffusion models, and diffusion posterior samplers, we find that pairing each initial noise with its negation consistently yields strongly negatively correlated samples. To explain this phenomenon, we combine experiments and theoretical analysis, leading to a symmetry conjecture that the learned score function is approximately affine antisymmetric (odd symmetry up to a constant shift), and provide evidence supporting it. Leveraging this negative correlation, we enable two applications: (1) enhancing image diversity in models like Stable Diffusion without quality loss, and (2) sharpening uncertainty quantification (e.g., up to 90% narrower confidence intervals) when estimating downstream statistics. Building on these gains, we extend the two-point pairing to a randomized quasi-Monte Carlo estimator, which further improves estimation accuracy. Our framework is training-free, model-agnostic, and adds no runtime overhead.
**footnotetext: Co-corresponding authors
@@ 1 Introduction
Diffusion models have set the state of the art in photorealistic image synthesis, high-fidelity audio, and video generation sohl2015deep ; ho2020denoising ; song2021scorebased ; kong2021diffwave ; they also power applications such as text-to-image generation ldm2022high , image editing and restoration meng2022sdedit , and inverse problem solving song2022solving .
@@
Much of the literature improves the first two ingredients and clusters into two strands: (i) architectural and training developments, which improve sample quality and scalability through backbone or objective redesign (e.g., EDM (karras2022elucidating, ), Latent Diffusion Models (ldm2022high, ), DiT (peebles2023scalable, )), and (ii) accelerated sampling, which reduces the number of denoising steps while retaining high-quality generation (e.g., DDIM (songj2021denoising, ), Consistency Models (song2023consistency, ), DPM-Solver++ (lu2023dpmsolver, ), and progressive distillation (salimans2022progressive, )).
However, the third ingredient—the initial Gaussian noise—has received comparatively little attention. Previous work examines it mostly through task-specific lenses: InitNO guo2024initno separates Gaussian noise seeds into “valid” and “invalid” groups according to their attention-score profiles. Qi et al. qi2024not show that inversion-stable noises reliably boost the quality of generated images. Zhou et al. zhou2024golden propose a learning framework to select high quality noises. The Crystal Ball detector ban2025the pinpoints “trigger patches” as crucial elements in object synthesis. FIND chen2024find optimizes the initial noise distribution so that outputs better match user prompts. CCS song2025ccs exploits an almost linear noise-to-image mapping to achieve controllable generation. Nevertheless, a systematic analysis of how the initial noise itself shapes the final outputs is still missing.
@@
Our perspective is orthogonal to prior work. Our first contribution is a simple yet model-agnostic finding: across unconditional, latent text-conditioned, and posterior-sampling diffusion models, pairing every Gaussian noise $z$ with its negation $-z$—known as antithetic sampling mcbook —consistently produces samples that are strongly negatively correlated. This phenomenon is universal as it holds regardless of architecture, dataset, sampling schedule, and both conditional and unconditional sampling. To illuminate this phenomenon, we conduct explanatory experiments and theoretical analysis, conjecture that the learned score function exhibits approximate affine symmetry, and present initial supporting evidence.
(a) Stable Diffusion
(b) CelebA-HQ
Figure 1: Antithetic noise produces visually “opposite” images. Left: a Stable Diffusion sketch with the prompt “cover, hyperdetailed painting, luminism, Bar lighting, complex, head and shoulders portrait.” Right: images from independent random-noise pairs (top) vs. an antithetic pair (bottom).
Leveraging this universal property, we present two concrete applications:
(i) Enhanced image diversity. As shown in Figure 1, negative correlation implies diversity. Because each antithetic noise pair drives the reverse-diffusion trajectories toward distant regions of the data manifold, the paired-sampling scheme markedly increases diversity “for free” compared with independent noise inputs, while still maintaining high image quality. The procedure is fully plug-and-play: one simply feeds the model with antithetic noise vectors.
(ii) Sharper uncertainty quantification. Because each antithetic pair generates strongly negatively correlated samples, these outputs form an effective control-variate pair. We apply this idea in an antithetic estimator for downstream statistics, reducing the 95% confidence-interval width by up to 90% and cutting computational effort by about 100 times compared to standard Monte Carlo sampling with independent noise.
@@
Building on the variance gains from antithetic pairs, we generalize the idea to apply quasi-Monte Carlo (QMC) and randomized QMC (RQMC). The resulting RQMC estimator often delivers further variance reduction. Although QMC has been widely used in computer graphics keller1995quasi ; waechter2011quasi , quantitative finance joy1996quasi ; l2009quasi , and Bayesian inference buchholz2018quasi ; liu2021quasi , this is, to our knowledge, its first application to diffusion models.
Together, these results suggest that principled manipulations in the initial-noise space can be a promising avenue for further gains in diffusion model’s efficiency, output diversity, and reliability.
@@
The remainder of the paper is organized as follows. Section 2 defines the problem and outlines our motivation. Section 3 presents our central finding that antithetic noise pairs produce strongly negatively correlated outputs. We offer both theoretical and empirical explanations for this phenomenon, and present the symmetry conjecture in Section 3.2. Section 4 details the construction of point estimates and confidence intervals using antithetic sampling for estimating downstream statistics, illustrates how antithetic sampling tightens confidence bounds, and discusses the generalization to QMC. Section 5 reports experiments on the aforementioned applications. Finally, Section 6 discusses our method’s potentials, limitations, and future directions. Appendix A, B, and C include proofs, additional experiments and detailed setups, and supplemental visualizations, respectively.
2 Setup and motivation
@@
A diffusion model aims to generate samples from an unknown data distribution $p_{0}$. It first noises data towards a standard Gaussian by applying progressively stronger Gaussian perturbations, then learns to reverse the process so that pure white noise can be denoised step-by-step back into target samples. The forward process simulates a stochastic differential equation (SDE): $\mathrm{d}\mathbf{x}_{t}=\mu(\mathbf{x}_{t},t)\mathrm{d}t+\sigma_{t}\mathrm{d}\mathbf{w}_{t},$ where $\{\mathbf{w}_{t}\}_{t=0}^{T}$ denotes the standard Brownion motion, and $\mu(\mathbf{x},t),\sigma_{t}$ are chosen by the users. Let $p_{t}$ denote the distribution of $\mathbf{x}_{t}$, where $\mathbf{x}_{0}\sim p_{0}$. A key result in song2021scorebased states that if we sample $\mathbf{y}_{T}\sim p_{T}$ and simulate the probability-flow ordinary differential equation (PF-ODE) backward from time $T$ to 00 as
@@
then for every time $t$ the marginal distribution of $\mathbf{y}_{t}$ coincides with $p_{t}$. Thus in an idealized world, one can perfectly sample from $p_{0}$ by simulating the PF-ODE (1).
In practice, the score function $\nabla\log p_{t}(\mathbf{x},t)$ is unavailable, and a neural network $\epsilon_{\theta}^{(t)}(\mathbf{x})$ is trained to approximate it, where $\theta$ represents its weights. The drift term $\mu$ and diffusion term $\sigma$ are designed such that $p_{t}$ converges to a standard Gaussian as $t\rightarrow\infty$. Therefore, one can generate new samples from $p_{0}$ by first sampling a Gaussian noise and simulating the PF-ODE (1) through some numerical integrator from $T$ to 00. For example, DDIM songj2021denoising has the (discretized) forward process as $\mathbf{x}_{k}\mid\mathbf{x}_{k-1}\sim\mathbb{N}(\sqrt{\alpha_{k}}\mathbf{x}_{k-1},(1-\alpha_{k})I)$ for $k=0,1,\ldots,T-1$, and backward sampling process as
@@
We remark that samples from $p_{0}$ can also be drawn by simulating the backward SDE with randomness at each step, as in the DDPM sampler ho2020denoising ; song2021scorebased . Throughout the main text, we focus on deterministic samplers such as DDIM to explain our idea. In Appendix B, we present additional experiments showing that our findings also extend to stochastic samplers like DDPM.
@@
In Stable Diffusion and its successors SDXL podell2024sdxl and SDXL Turbo sauer2024adversarial , a pretrained VAE first compresses each image to a latent tensor $z\in\mathcal{Z}$. A frozen text encoder embeds the prompt as $c\in\mathbb{R}^{m}$. During training, the network $\epsilon_{\theta}^{(t)}$ receives $(z_{t},c)$ and learns $\nabla_{z_{t}}\log p_{t}(z_{t}\mid c)$. At generation time we draw latent noise $z_{T}\sim\mathbb{N}(0,I)$ and run the reverse diffusion from $t=T$ to 00, yielding a denoised latent $z_{0}$, which the decoder maps back to pixels. Thus, given a prompt $c$ and an initial Gaussian noise, the sampler produces an image from $p_{0}(\cdot\mid c)$.
@@
Diffusion model has also been utilized as a prior for posterior sampling in many inverse problems. Suppose we observe only a partial or corrupted measurement $\mathbf{y}_{\text{obs}}=\mathcal{A}(\mathbf{x})+\text{noise}$ from an unknown signal $\mathbf{x}$ via a known forward operator $\mathcal{A}$. From a Bayesian perspective, we impose a prior $p(\mathbf{x})$ on $\mathbf{x}$, and perform inference based on the posterior $p(\mathbf{x}\mid\mathbf{y}_{\text{obs}})\;\propto\;p(\mathbf{y}_{\text{obs}}\mid\mathbf{x})\,p(\mathbf{x}).$ Diffusion posterior sampling arises when $p(\mathbf{x})$ is given by a pretrained diffusion model. Given initial noise $z$ and observed $\mathbf{y}_{\text{obs}}$, diffusion posterior samplers apply a sequence of denoising steps from $T$ to 00, each step resembling (2) but incorporating $\mathbf{y}_{\text{obs}}$  chung2023diffusion ; song2022solving ; song2024solving .
@@ 2.1 Motivation
Beyond examining how variations in the initial noise affect the outputs, several technical considerations motivate our focus on the antithetic noise pair $(z,-z)$ for diffusion models. Let DM denote certain diffusion model’s mapping from an initial noise vector to a generated sample.
[…]
  • Maximal separation in the noise space: High-dimensional Gaussians concentrate on the sphere of radius $\sqrt{\text{dimension}}$ vershynin2018high , so any draw $z$ and its negation $-z$ lie at opposite poles of that hypersphere. This antipodal pairing represents the maximum possible perturbation in noise space, making it a natural extreme test of the sampler’s behavior.
  • Measuring (non-)linearity: Recent work has examined how closely score networks approximate linear operators. Empirical studies across various diffusion settings show that these networks behave as locally linear maps; this behavior forms the basis for new controllable sampling schemes chen2024exploring ; li2024understanding ; song2025ccs . Since linear transformations preserve correlation, we can use correlation to evaluate linearity: if a score network were exactly linear, feeding it with noises $z$ and $-z$ would yield a correlation of $-1$ between $\text{DM}(z)$ and $\text{DM}(-z)$. Thus, the difference between the observed correlation and $-1$ provides a direct measure of the network’s departure from linearity.
@@ 3 Negative correlation from antithetic noise
3.1 Pixel-wise correlations: antithetic vs. independent noise
Table 1 reports Pearson correlations between pixel values of image pairs produced by pretrained DDIM models under two noise schemes—PN (positive vs. negative: $z$ vs. $-z$) and RR (random vs. random: $z_{1}$ vs. $z_{2}$)—across CIFAR-10 cifar-10 , CelebA-HQ celeba-hq , Church yu2015lsun , and Stable Diffusion.
  • CIFAR-10, CelebA-HQ, Church: For each of these three datasets, we draw $1600$ antithetic noise pairs, compute the pixelwise correlation for each generated image pair, and report their average. We repeat the same procedure with $1600$ independent random noise pairs for comparison.
  • Stable Diffusion: Since this is a prompt-based model, we use two common benchmarks (DrawBench and Pick-a-Pic) and randomly select 100 prompts from each. For each prompt, we generate $100$ antithetic and $100$ independent noise pairs, measuring and averaging the resulting correlations.
Across all datasets, PN pairs have markedly and consistently lower pixel correlations than RR pairs. Their average correlation never surpasses $-0.33$ and can dip below $-0.7$. In contrast, RR pairs show mildly positive correlations, reflecting the consistent overall outlines enforced by the model’s architecture and training dataset. Moreover, the 75% quantile of PN correlation is frequently even lower than the lowest RR correlation. The histogram of pixel-wise correlation coefficients is plotted in Figure 2. The PN pair correlations are heavily left-skewed—most coefficients range from $-0.9$ to $-0.4$—whereas the RR correlations span a broader, nearly symmetric range centered just above zero.
@@
The same behavior appears in both DDPM models and diffusion posterior samplers, which we report in Appendix B.1. These results demonstrate that the negative correlation resulted from antithetic sampling is a universal phenomenon across various datasets and diffusion models.
Table 1: Pearson correlation coefficients for PN and RR pairs.
Statistic CIFAR10 CelebA Church Stable Diffusion
PN RR PN RR PN RR PN RR
Mean -0.76 0.05 -0.33 0.24 -0.62 0.08 -0.47 0.05
Min -0.97 -0.68 -0.87 -0.66 -0.95 -0.46 -0.71 -0.23
25th percentile -0.85 -0.11 -0.47 0.12 -0.70 -0.04 -0.54 -0.02
75th percentile -0.71 0.21 -0.22 0.39 -0.55 0.19 -0.40 0.11
Max 0.08 0.74 0.49 0.78 -0.21 0.67 -0.18 0.33
(a) CIFAR10
(b) LSUN-Church
(c) CelebA-HQ
Figure 2: Pearson Correlation histograms for PN and RR pairs across three datasets using DDIM. Dashed lines indicate the average Pearson correlation for each group.

### v2

@@ Abstract
We systematically study antithetic initial noise in diffusion models, discovering that pairing each noise sample with its negation consistently produces strong negative correlation. This universal phenomenon holds across datasets, model architectures, conditional and unconditional sampling, and even other generative models such as VAEs and Normalizing Flows. To explain it, we combine experiments and theory and propose a symmetry conjecture that the learned score function is approximately affine antisymmetric (odd symmetry up to a constant shift), supported by empirical evidence. This negative correlation leads to substantially more reliable uncertainty quantification with up to $90\%$ narrower confidence intervals. We demonstrate these gains on tasks including estimating pixel-wise statistics and evaluating diffusion inverse solvers. We also provide extensions with randomized quasi-Monte Carlo noise designs for uncertainty quantification, and explore additional applications of the antithetic noise design to improve image editing and generation diversity. Our framework is training-free, model-agnostic, and adds no runtime overhead. Code is available at https://github.com/jjia131/Antithetic-Noise-in-Diffusion-Models-page.
**footnotetext: Co‐corresponding authors
@@ 1 Introduction
Diffusion models have set the state of the art in photorealistic image synthesis, high‑fidelity audio, and video generation  (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2021b; Kong et al., 2021); they also power applications such as text-to-image generation  (Rombach et al., 2022), image editing and restoration  (Meng et al., 2022), and inverse problem solving  (Song et al., 2022).
@@
Much of the literature improves the first two ingredients and clusters into two strands: (i) architectural and training developments, which improve sample quality and scalability through backbone or objective redesign (e.g., EDM  (Karras et al., 2022), Latent Diffusion Models  (Rombach et al., 2022), DiT  (Peebles and Xie, 2023)), (ii) accelerated sampling, which reduces the number of denoising steps while retaining high-quality generation (e.g., DDIM  (Song et al., 2021a), Consistency Models  (Song et al., 2023), DPM‑Solver++  (Lu et al., 2023), and progressive distillation  (Salimans and Ho, 2022)).
However, the third ingredient—the initial Gaussian noise—has received comparatively little attention. Prior work has optimized initial noise for generation quality, editing, controllability, or inverse problem solving (Guo et al., 2024; Qi et al., 2024; Zhou et al., 2024; Ban et al., 2025; Chen et al., 2024a; Eyring et al., 2024; Song et al., 2025; Wang et al., 2024; Chihaoui et al., 2024). However, most of these efforts are task-specific. A systematic understanding of how noise itself shapes diffusion model outputs is still missing.
@@
Our perspective is orthogonal to prior work. Our central discovery is both simple and universal: pairing every Gaussian noise $z$ with its negation z-z—known as antithetic sampling (Owen, 2013)—consistently produces samples that are strongly negatively correlated. This phenomenon holds regardless of architecture, dataset, sampling schedule, and both conditional and unconditional sampling. It further extends to other generative models such as VAEs and Normalizing Flows. We explain this phenomenon through both experiments and theory. This leads to a symmetry conjecture that the score function is approximately affine antisymmetric, providing a new structural insight supported by empirical evidence.
Figure 1: Use antithetic noise z-z and $z$ (with condition $c$) to generate visually “opposite” images.
This universal property has direct impact on uncertainty quantification and also enables additional applications:
(i) Sharper uncertainty quantification. Antithetic pairs naturally act as control variates, enabling significant variance reduction and thus sharper uncertainty quantification. Our antithetic estimator delivers up to $90\%$ tighter confidence intervals and cuts computation cost by more than $100$ times. The efficiency gain immediately leads to huge cost savings in a variety of tasks, including bias detection in generation and diffusion inverse solver evaluation, as we demonstrate in later experiments.
(ii) Other applications. Because each antithetic noise pair drives reverse-diffusion trajectories toward distant regions of the data manifold (see Figure 1), the paired-sampling scheme increases diversity “for free” while preserving high image quality, as confirmed by SSIM and LPIPS in our experiments. Moreover, algorithms that rely on intermediate sampling or approximation steps also benefit from the improved reliability provided by antithetic noise. As an illustration, we present an image editing example in B.2 and show that our antithetic design serves as a plug-and-play tool to improve performance at no additional cost.
@@
Building on the antithetic pairs, we generalize the idea to apply quasi-Monte Carlo (QMC) and randomized QMC (RQMC). The resulting RQMC estimator often delivers further variance reduction. Although QMC has been widely used in computer graphics (Keller, 1995; Waechter and Keller, 2011), quantitative finance (Joy et al., 1996; L’Ecuyer, 2009), and Bayesian inference (Buchholz et al., 2018; Liu and Owen, 2021), this is, to our knowledge, its first application to diffusion models.
In summary, we discover a universal property of initial noise, reveal a new symmetry in score networks, and demonstrate concrete benefits in various practical applications. This positions initial noise manipulations as a simple, training-free tool for advancing generative modeling.
@@
The remainder of the paper is organized as follows. Section 2 defines the problem and outlines our motivation. Section 3 presents our central finding that antithetic noise pairs produce strongly negatively correlated outputs. We offer both theoretical and empirical explanations for this phenomenon, and present the symmetry conjecture in Section 3.2. Section 4 develops estimators and their confidence intervals via antithetic sampling, and extends the approach to QMC. Section 5 reports experiments on the aforementioned applications. Section 6 discusses our method and outlines future directions. Appendix A, B, and C include proofs, additional experiments and detailed setups, and supplemental visualizations, respectively.
2 Setup, motivation, and related works
@@
A diffusion model aims to generate samples from an unknown data distribution $p_{0}$. It first noises data towards a standard Gaussian progressively, then learns to reverse the process so that Gaussian noise can be denoised step-by-step back into target samples. The forward process simulates a stochastic differential equation (SDE): $\mathrm{d}\mathbf{x}_{t}=\mu(\mathbf{x}_{t},t)\mathrm{d}t+\sigma_{t}\mathrm{d}\mathbf{w}_{t},$ where $\{\mathbf{w}_{t}\}_{t=0}^{T}$ denotes the standard Brownian motion, and $\mu(\mathbf{x},t),\sigma_{t}$ are chosen by the users. Let $p_{t}$ denote the distribution of $\mathbf{x}_{t}$. Song et al. (2021b) states that if we sample $\mathbf{y}_{T}\sim p_{T}$ and simulate the probability-flow ordinary differential equation (PF-ODE) backward from time $T$ to 0 as
@@
then for every time $t$ the marginal distribution of $\mathbf{y}_{t}$ coincides with $p_{t}$. Thus, in an idealized world, one can perfectly sample from $p_{0}$ by simulating the PF-ODE (1).
In practice, the score function $\nabla\log p_{t}(\mathbf{x},t)$ is unavailable, and a neural network $\epsilon_{\theta}^{(t)}(\mathbf{x})$ is trained to approximate it, where $\theta$ denotes its weights. Therefore, one can generate new samples from $p_{0}$ by first sampling a Gaussian noise and simulating the PF-ODE (1) through a numerical integrator from $T$ to 0. For example, DDIM (Song et al., 2021a) has the (discretized) forward process as $\mathbf{x}_{k}\mid\mathbf{x}_{k-1}\sim\mathbb{N}(\sqrt{\alpha_{k}}\mathbf{x}_{k-1},(1-\alpha_{k})I)$ for $k=0,\ldots,T-1$, and backward sampling process as
@@
We remark that samples from $p_{0}$ can also be drawn by simulating the backward SDE with randomness at each step, as in the DDPM sampler (Ho et al., 2020; Song et al., 2021b). Throughout the main text, we focus on deterministic samplers such as DDIM to explain our idea. In Appendix B.4, we present additional experiments showing that our findings also extend to stochastic samplers like DDPM.
@@
In Stable Diffusion and its successors SDXL and SDXL Turbo (Podell et al., 2024; Sauer et al., 2024), a pretrained VAE first compresses each image to a latent tensor $z$. A text encoder embeds the prompt as $c\in\mathbb{R}^{m}$. During training, the network $\epsilon_{\theta}^{(t)}$ receives $(z_{t},c)$ and learns $\nabla_{z_{t}}\log p_{t}(z_{t}\mid c)$. At generation time we draw latent noise $z_{T}\sim\mathbb{N}(0,I)$ and run the reverse diffusion from $t=T$ to 0, yielding a denoised latent $z_{0}$, which the decoder maps back to pixels. Given a prompt $c$ and an initial Gaussian noise, the sampler produces an image from $p_{0}(\cdot\mid c)$.
@@
Diffusion model can also be used as a prior in inverse problems. Suppose we observe only a partial or corrupted measurement $\mathbf{y}_{\text{obs}}=\mathcal{A}(\mathbf{x})+\text{noise}$ from an unknown signal $\mathbf{x}$. Diffusion posterior sampling aims to sample from the posterior distribution $p(\mathbf{x}\mid\mathbf{y}_{\text{obs}})\;\propto\;p(\mathbf{y}_{\text{obs}}\mid\mathbf{x})\,p(\mathbf{x})$, where $p(\mathbf{x})$ is given by a pretrained diffusion model. Given initial noise $z$ and observed $\mathbf{y}_{\text{obs}}$, diffusion posterior samplers apply a sequence of denoising steps from $T$ to 0, each step resembling (2) but incorporating $\mathbf{y}_{\text{obs}}$ (Chung et al., 2023a; Song et al., 2022, 2024).
@@ 2.1 Motivation
Beyond examining how variations in the initial noise affect the outputs, several technical considerations motivate our focus on the antithetic noise pair (z,z)(z,-z) for diffusion models. Let DM denote certain diffusion model’s mapping from an initial noise vector to a generated image sample.
  • Preserving quality: Since $z\sim\mathbb{N}(0,I)$ implies $z\sim\mathbb{N}(0,I)$, the initial noises $z$ and z-z share the same marginal distribution. Consequently, $\text{DM}(z)\overset{d}{=}\text{DM}(-z)$, so all per‐sample statistics remain unchanged. In other words, negating the initial noise does not degrade generation quality.
  • Maximal separation in the noise space: High-dimensional Gaussians concentrate on the sphere of radius $\sqrt{\text{dimension}}$ (Vershynin, 2018), so any draw $z$ and its negation z-z lie at opposite poles of that hypersphere. This antipodal pairing represents the maximum possible perturbation in noise space, making it a natural extreme test of the sampler’s behavior.
  • Measuring (non-)linearity: Recent work has examined how closely score networks approximate linear operators. Empirical studies across various diffusion settings show that these networks behave as locally linear maps; this behavior forms the basis for new controllable sampling schemes (Chen et al., 2024b; Li et al., 2024b; Song et al., 2025). Correlation is a natural choice to measure linearity: if a score network were exactly linear, feeding it with noises $z$ and z-z would yield a correlation of 1-1 between $\text{DM}(z)$ and $\text{DM}(-z)$. Thus, the difference between the observed correlation and 1-1 provides a direct measure of the network’s departure from linearity.
@@ 3 Negative correlation from antithetic noise
3.1 Pixel‐wise correlations: antithetic vs. independent noise
We compare the similarity between paired images generated under two sampling schemes: PN (positive vs. negative, $z$ vs. z-z) and RR (random vs. random, $z_{1}$ vs. $z_{2}$) under three settings: (1) unconditional diffusion models, (2) class- or prompt-based conditional diffusion models, and (3) generative models beyond diffusion.
For diffusion models, we evaluate both unconditional and conditional generation using publicly available pre-trained checkpoints. Notably, we include both traditional U-Net architecture and transformer-based DiT  (Peebles and Xie, 2023). Implementation details are described in Section 5.1. We also test on distilled diffusion models  (Song et al., 2023) using consistency distillation checkpoints. For generative models beyond diffusion, we select two representative baselines: an unconditional VAE on MNIST and a conditional Glow model  (Kingma and Dhariwal, 2018) on CIFAR-10. The experimental details for consistency models, VAE and Glow are provided in Appendix B.1.
To quantify similarity, we use two metrics: the standard Pearson correlation and a centralized Pearson correlation. Let $x_{i,1}$ and $x_{i,2}$ denote the flattened pixel values of the two images in the $i$-th generated pair. The standard Pearson correlation is computed directly between $x_{i,1}$ and $x_{i,2}$. To correct for dataset-level or class-level bias, we also define a centralized correlation. For a dataset, prompt, or class with $K$ pairs generated, we compute the mean $\mu_{c}=\sum_{i=1}^{K}(x_{i,1}+x_{i,2})/2K$. The centralized correlation of the $i$-th pair is defined as the standard Pearson correlation between the centralized images $x_{i,1}-\mu_{c}$ and $x_{i,2}-\mu_{c}$. For each comparison, $t$-test is conducted to assess statistical significance. In all experiments, the resulting $p$-values are negligible ($10^{-10}$), which confirms significance; hence, they are omitted from the presentation.
Additional metrics such as the Wasserstein distance are presented in Appendix B.4.
Results: Table 1 summarizes the statistics in all classes of models. PN pairs consistently show significantly stronger negative correlations than RR pairs, with this contrast also visually evident in their histograms (Figure 2). In addition, centralization strengthens the negative correlation, since it removes shared patterns. For example, in CelebA-HQ, centralization removes global facial structure, while in DiT and Glow, it removes class-specific patterns.
@@
The same behavior appears in both DDPM models and diffusion posterior samplers, which we report in Appendix B.4. These results demonstrate that the negative correlation resulting from antithetic sampling is a universal phenomenon across diverse architectures and conditioning schemes.
Table 1: Correlation results across different models and datasets, shown are means (SD). Rows 1–3 are pretrained unconditional diffusion models on different datasets. Rows 4–5 are conditional diffusion models. Rows 6–8 are pretrained consistency models on different datasets. Rows 9–10 are generative models that are not diffusion-based.
Model Dataset Standard Correlation Centralized Correlation
PN $R$ PN $R$
Uncon. Model LSUN-Church -0.62 (0.11) 0.08 (0.17) -0.77 (0.07) 0.00 (0.17)
CelebA-HQ -0.34 (0.19) 0.25 (0.20) -0.78 (0.06) -0.01 (0.20)
CIFAR-10 -0.76 (0.13) 0.05 (0.24) -0.86 (0.07) 0.00 (0.23)
SD 1.5 DiT LAION-2B(en) -0.47 (0.05) 0.05 (0.04) -0.54 (0.08) 0.00 (0.01)
ImageNet-256 -0.07 (0.27) 0.26 (0.18) -0.45 (0.08) 0.00 (0.03)
Consistency Model LSUN-Cat -0.88 (0.06) 0.02 (0.14) -0.91 (0.05) 0.01 (0.14)
LSUN-Bedroom -0.78 (0.07) 0.03 (0.15) -0.84 (0.06) 0.00 (0.15)
ImageNet-64 -0.71 (0.14) 0.02 (0.19) -0.75 (0.12) 0.01 (0.19)
VAE Glow MNIST 0.21 (0.12) 0.42 (0.17) -0.41 (0.12) -0.00 (0.24)
CIFAR-10 -0.52 (0.02) 0.08 (0.05) -0.57 (0.02) -0.01 (0.02)
(a) CelebA-HQ
(b) DiT (class 919)
(c) Glow (class 0)
Figure 2: Histograms of standard and centralized Pearson correlation coefficients for CelebA-HQ, and DiT class and Glow in single class. Dashed lines indicate the average.